OpenStudy (anonymous):

If f'(x) = 2/x and f(sqrt of e) =5, then f(e) =?

OpenStudy (anonymous):

\[f(\sqrt{e})=5\]??

OpenStudy (anonymous):

are you allowed to integrate? if so, then if \[f'(x)=\frac{2}{x}\] then \[f(x)=2\ln(x)+C\]

OpenStudy (anonymous):

we can find C easily enough, since \[f(e)=5=2\ln(e)+C=2+C\] making \(C=3\) and so your funciton is \[f(x)=2\ln(x)+3\]

OpenStudy (anonymous):

Then u plug in e right? But u get 2ln (e) +3 which is 5 but in teh answer key it says its 6

OpenStudy (anonymous):

oh because i am an idiot

OpenStudy (anonymous):

it says \[f(\sqrt{e})=5\]

OpenStudy (anonymous):

so we have \[f(x)=2\ln(x)+C\] and we know that \[f(\sqrt{e})+C=5\] so \[2\ln(\sqrt{e})+C=5\] \[1+C=5\] and so \[C=4\]

OpenStudy (anonymous):

therefore your function is \[f(x)=2\ln(x)+4\] and then of course \[f(e)=2\ln(e)+4=2+4=6\] sorry about my original mistake

OpenStudy (anonymous):

OH no problem! Thanks for finxing it! =)